OpenStudy (anonymous):

please help!!! ill give a medal to anyone that help im so lost on what to do solve for x: 2/5(x-2)=4x

OpenStudy (anonymous):

2/5 * (x - 2) = 4x [2(x - 2)]/5 = 4x 2(x) - 2(2) = 5(4x) 2x - 4 = 20x 2x - 20x = 4 -18x = 4 x = 4/(-18) x = -2/9

OpenStudy (anonymous):

okay i think i understand thanks